Assume that agriculture prices fall and the farming sector faces a mild recession. The demand for the small tractors drops to: P = 26,000 – Q. 

Suppose the recession is only temporary, and demand will recover soon. What output adjustment should the firm make during the recession?


1
Expert's answer
2022-02-20T16:27:28-0500

New demand function is:

"P= 26000-Q"

The firm maximazes output when "MR=MC"

"MR=26000-2Q"

"MC=20000"

"26000-2Q=20000"

"2Q=6000"

"Q=3000"